# Positive mass and isoperimetry for continuous metrics with nonnegative scalar curvature

### v2

@@ 1.4. The question of rigidity
In the smooth global PMT [SchoenYauPMT], one gets that the mass is zero if and only if the metric is flat. Restricting the attention to the smooth case, an easy combination of the local IMCF arguments leading to Theorem 1.2 with the rigidity triggered by the constancy of the Hawking mass along the weak IMCF [HuiskenIlmanen] allows us to prove the following quasi-local counterpart.
@@
for all the domains $\partial E_{t}\Subset\Omega$ evolving through the local IMCF, where $\mathfrak{m}_{H}(\partial E_{t})$ denotes the Hawking mass (3.25) of their boundary. We can then easily conclude invoking a rigidity argument by Huisken–Ilmanen [HuiskenIlmanen]. We observe that a counterpart of Theorem 1.4, with (1.5) replaced by
@@
has been already obtained by Mondino–Templeton-Browne by different means [t.browne-mondino]. In fact, we stress that having proved the existence of a weak local IMCF allows us to derive a different proof of the result in [t.browne-mondino]. Indeed, (1.7) would directly imply (1.6), thus allowing us to conclude using the rigidity argument of Huisken–Ilmanen as mentioned above, see also the proof of (1.5) below. Notice that the condition (1.5) makes sense also for $C^{0}$-metrics.
@@
Acknowledgments. A part of this research was carried out while G.A., M. F., and M.P. were visiting S.N. at UFABC. The joyful working conditions and the hospitality are gratefully acknowledged. Another part of the work was carried out while G. A. and M. F. were participating in the workshop “Recent Advances in Comparison Geometry” in Hangzhou. They thank the Banff International Research Station and the Institute of Advanced Studies in Mathematics for this enriching opportunity. G.A. is grateful to Paula Burkhardt-Guim for useful discussions around the topic of the paper. M. F. is grateful to Andrea Mondino for his interest and for having brought to his attention the paper [t.browne-mondino]. M.P. is grateful to Roberto Ladu for clarifying discussions and stimulating advice. The authors are grateful to Luca Benatti, Luciano Mari, and Kai Xu for useful discussions around the topic of this paper. They are also grateful to Virginia Agostiniani, Alessandro Carlotto, Jeff Jauregui, Dan Lee, Lorenzo Mazzieri, and Christina Sormani for their interest in this work.
@@ 2.1. Basic definitions and properties of $C^{0}$-metrics
We briefly recall some metric properties of $C^{0}$-Riemannian manifolds $(M,g)$, and we refer the reader to [burtscher-continuousmetrics, petrunin-metricgeometrytwolectures] for a more detailed discussion. As in the smooth case, the continuous Riemannian metric $g$ defines a length structure on absolutely continuous curves on $M$, which in turn gives raise to a distance $\mathsf{d}$ that induces the manifold topology. We always assume that $(M,\mathsf{d})$ is a complete metric space, so that by Hopf–Rinow Theorem the distance $\mathsf{d}$ is geodesic, and every closed ball is compact.
@@
the map $F|_{B_{R}(o)}:(B_{R}(o),\mathsf{d})\to(F(B_{R}(o)),\tilde{\mathsf{d}})$ is $(1+\delta)$-biLipschitz with its image.
@@
[…]
In the notation of Section 2.1, arguing as in the proof of Section 2.1, it follows that for any $R,\varepsilon>0$ the map $F_{i}:(B_{R}(p),\mathsf{d})\to(F_{i}(B_{R}(p)),\mathsf{d}_{i})$ is $(1+\varepsilon)$-biLipschitz with its image for any $i$ large enough.
[…]
@@
Concerning the second part of the statement, let $r>0$ to be chosen small and fix a local chart $\varphi:B_{10r}(x_{0})\to\varphi(B_{10r}(x_{0}))=:\Omega\subset\mathbb{R}^{n}$ such that $\varphi(x_{0})=0$, and denote $g_{0}:=g_{\rm eu}$. We identify $B_{10r}(x_{0})$ with $(\Omega,g)$ in the local chart. Let $\tilde{\mathsf{d}}$ be as in Section 2.1 with $N=\mathbb{R}^{n}$. By continuity of the metric tensor, for any $\varepsilon>0$ we can take $r$ so small that $|g_{x}(v,v)-g_{0}(v,v)|\leq\varepsilon g_{0}(v,v)$ for any $x\in\Omega$ and any $v\in\mathbb{R}^{n}$. For $\varepsilon$ small enough, it follows from Section 2.1 that the identity ${\rm id}:(B^{g}_{r}(0),\mathsf{d})\to(B^{g}_{r}(0),\tilde{\mathsf{d}})$ is $(1+\delta)$-biLipschitz. It remains to observe that $\tilde{\mathsf{d}}=\mathsf{d}_{\rm eu}$ on $B^{g}_{r}(0)$. Indeed, fix $p,q\in B^{g}_{r}(0)$ and let $\gamma_{i}:[0,1]\to\Omega$ be a sequence of constant speed curves such that $\int_{0}^{1}|\gamma_{i}^{\prime}|_{\rm eu}\to\tilde{\mathsf{d}}(p,q)$. It follows that $\gamma_{i}([0,1])\subset B^{g}_{9r}(0)$ for large $i$. Indeed, taking $\sigma:[0,1]\to B^{g}_{5r}(0)$ a constant speed geodesic for $\mathsf{d}$ from $p$ to $q$, we know that
@@ 2.2. BV functions and sets of finite perimeter
Let $(M,g)$ be a complete $C^{0}$-Riemannian manifold of dimension $n$. Then we can consider the complete and separable metric measure space $(M,\mathsf{d},\mathcal{H}^{n})$. Following [MIRANDA2003, ADM14], we define the total variation $|Df|(B)\in[0,+\infty]$ of a function $f\in L^{1}_{\mathrm{loc}}(M)$ in a Borel set $B\subseteq M$ as
@@
then plugging ${\rm lip\,}f_{n}$ in place of $|\nabla f_{n}|$ in (2.1) yields the same quantity. Indeed, if $f\in\operatorname{Lip}_{\rm loc}(M)$, then ${\rm lip\,}f=|\nabla f|$ almost everywhere, see Appendix A.
@@
Given any $f\in\mathrm{Lip}_{\mathrm{loc}}(M)$, it holds that $|Df|$ is a locally finite measure and $|Df|={\rm lip\,}f\mathcal{H}^{n}=|\nabla f|\mathcal{H}^{n}$, see Appendix A. Moreover, we recall the coarea formula in our setting,
Theorem 2.1 (Coarea formula [MIRANDA2003, Proposition 4.2]).
@@ 2.3. Technical Lemmas on $C^{0}$-Riemannian manifolds
In this section we prove several technical lemmas about $C^{0}$-Riemannian manifolds we are going to use throughout the paper. Notably, we prove: a precompactness theorem for $BV$ functions on converging sequences of continuous Riemannian manifolds (Section 2.3); the fact that every continuous Riemannian manifold is PI on every ball (Section 2.3); continuity of the isoperimetric profile of continuous Riemannian manifolds with $C^{0}$-controlled geometry at infinity (Section 2.3); and a generalized existence theorem for the isoperimetric problem on continuous Riemannian manifolds with $C^{0}$-controlled geometry at infinity (Theorem 2.2).
[…]
@@
For any $\varepsilon\in(0,1)$ let $g^{\varepsilon}$ be a smooth Riemannian metric on $\overline{B}_{20r}^{g}(p)$ such that $|g(v,v)-g^{\varepsilon}(v,v)|\leq\varepsilon g(v,v)$ for any $v\in T_{x}M$ and $x\in B_{20r}(p)$. Gluing the boundary of a smooth connected open domain $D$ such that $\overline{B}_{10r}^{g}(p)\subset D\subset\subset B_{20r}^{g}(p)$ with a half-cylinder $[0,+\infty)\times\partial D$ and suitably extending the metric $g^{\varepsilon}$, we obtain a smooth complete Riemannian manifold $(N,g^{\varepsilon})$. Denote by $\mathsf{d}^{\varepsilon}$ the Riemannian distance on $(N,g^{\varepsilon})$ and by $\tilde{\mathsf{d}}$ the distance defined in Section 2.1 with $F=\iota:B_{10r}^{g}(p)\to B_{10r}^{g}(p)\subset(N,g^{\varepsilon})$. Arguing as in the second part of the proof of Section 2.1, it follows that $\tilde{\mathsf{d}}=\mathsf{d}^{\varepsilon}$ on $B_{r}^{g}(p)$. Indeed, as in the proof of Section 2.1 we find that for any $x,y\in B^{g}_{r}(p)$ there exists a constant speed curve $\gamma:[0,1]\to B^{g}_{10r}(p)$ such that $\tilde{\mathsf{d}}(x,y)=\int_{0}^{1}|\gamma^{\prime}|_{g^{\varepsilon}}$, hence $\gamma$ is a geodesic for the metric $g^{\varepsilon}$. For every $x\in B^{g}_{r}(p)$, we have that for almost every $y\in B^{g}_{r}(p)$ there exists a unique, hence minimizing, geodesic in $(N,g^{\varepsilon})$ from $x$ to $y$. Then, for every $x\in B^{g}_{r}(p)$, we have that for almost every $y\in B^{g}_{r}(p)$ the curve $\gamma$ just obtained must be the minimizing geodesic from $x$ to $y$ in $(N,g^{\varepsilon})$. Hence, given $x\in B^{g}_{r}(p)$, there holds $\tilde{\mathsf{d}}(x,y)=\mathsf{d}^{\varepsilon}(x,y)$ for almost every $y\in B^{g}_{r}(p)$. Thus by continuity $\tilde{\mathsf{d}}=\mathsf{d}^{\varepsilon}$ on $B^{g}_{r}(p)$. Therefore choosing $\varepsilon$ small enough and eventually renaming $g^{\varepsilon}$ into $g^{\delta}$, the conclusion follows from Section 2.1. ∎
[…]
@@
Indeed, fix on a compact ball $B\subset M$ a sequence $g_{i}$ of smooth metrics such that $g_{i}\to g$ uniformly on $B$ (see, e.g., Section 2.3). Then, up to subsequences, $\partial^{e}E\cap B$ does not depend on the metrics $g_{i},g$ chosen in the definition (2.3) (compare also with item (1) in Section 2.3). Moreover, by the very definition of Hausdorff measure, $(1-o_{i}(1))\mathcal{H}^{n-1}_{g_{i}}\llcorner B\leq\mathcal{H}^{n-1}_{g}\llcorner B\leq(1+o_{i}(1))\mathcal{H}^{n-1}_{g_{i}}\llcorner B$. From the classical De Giorgi-Federer’s theorem, $|D\chi_{E}|_{i}(\Omega)=\mathcal{H}^{n-1}_{g_{i}}(\partial^{e}E\cap\Omega)$ for every $\Omega\subset\subset B$. Thus, taking into account Appendix A one has that $|D\chi_{E}|_{i}(\Omega)\to|D\chi_{E}|(\Omega)$ for every Borel $\Omega\subset\subset B$, and thus we conclude $|D\chi_{E}|\llcorner B=\mathcal{H}^{n-1}\llcorner(\partial^{e}E\cap B)$. Since $B$ was arbitrary, we get the sought claim. In the following, if $E$ is a set of locally finite perimeter, integrals over its essential boundary are tacitly understood to be taken with respect to the perimeter measure.
[…]
Let $(M_{i},g_{i},p_{i})$ be a sequence of pointed $C^{0}$-Riemannian manifolds of dimension $n$ converging in $C^{0}$-sense to a pointed $C^{0}$-manifold $(M,g,p)$. Denote by $\mathsf{d},\mathsf{d}_{i}$ the Riemannian distances on $(M,g),(M_{i},g_{i})$ respectively. Let $f_{i}\in BV(M_{i})$ be such that $\sup_{i}\|f_{i}\|_{L^{1}(M_{i})}+|Df_{i}|(M_{i})<+\infty$. Then, up to subsequence, there exist $f\in BV(M)$, $R_{i}\nearrow+\infty$ and $(1+1/i)$-biLipschitz embeddings $F_{i}:(B_{R_{i}}(p),\mathsf{d})\to(M_{i},\mathsf{d}_{i})$ with $F_{i}(p)=p_{i}$ such that the functions $f_{i}\circ F_{i}$ converge to $f$ in $L^{1}_{\rm loc}(M)$. Moreover
@@
By a diagonal argument, up to passing to a subsequence, by Section 2.1 there exist $R_{i}\nearrow+\infty$ and $(1+1/i)$-biLipschitz embeddings $F_{i}:B_{R_{i}}(p)\to M_{i}$ with $F_{i}(p)=p_{i}$. Denote $h_{i}\vcentcolon=f_{i}\circ F_{i}$. For any $r>0$, for $i$ large we have that $h_{i}\in BV(B_{r}(p))$ with $\sup_{i}\|h_{i}\|_{L^{1}(B_{r}(p))}+|Dh_{i}|(B_{r}(p))<+\infty$. If we show that $h_{i}$ admits a subsequence converging in $L^{1}(B_{r/2}(p))$, the first part of the statement follows. Indeed, by Section 2.3 we can find a smooth Riemannian manifold $(N,g^{\delta})$ such that it is well defined the inclusion $\iota:(B_{r}(p),\mathsf{d})\to(N,\mathsf{d}^{\delta})$ and $\iota$ is $2$-biLipschitz with its image. Hence $h_{i}$ can be seen as an equibounded sequence in $BV(B_{r}^{g}(p),g^{\delta})$. By classical precompactness we can extract a subsequence converging in $L^{1}(B_{r/2}^{g}(p),\mathcal{H}^{n}_{g^{\delta}})$. Since $\mathcal{H}^{n}_{g^{\delta}}$ and $\mathcal{H}^{n}_{g}$ are equivalent, the subsequence converges in $L^{1}(B_{r/2}(p))$ as well.
@@
The claims are well-known to hold true on manifolds with smooth Riemannian metrics as a consequence of the fact that one can always find a lower bound for the Ricci curvature on compact sets. In fact, on a smooth manifold, (1) is a consequence of the Bishop–Gromov comparison theorem [PetersenBook, Lemma 7.1.4], (2) follows from [Buser, RajalaPoincare], and (3) is a consequence of (1) and (2) by [HajlaszKoskela, Theorem 9.7] and [Ambrosio02, Theorem 4.3, Remark 4.4]. Therefore claims (1) and (3) in the setting of the statement readily follow from Section 2.3 applied with $r$ sufficiently large.
It remains to prove (2). Let $(N,g^{\delta})$ be given by Section 2.3 with $\delta=1/4$ and radius equal to $3R$. Letting now $f\in\operatorname{Lip}_{\rm loc}(M)$, for $x\in B_{R}(p)$ and $r\leq R$, we have $h\vcentcolon=f\circ\iota^{-1}\in\operatorname{Lip}_{\rm loc}(\iota(B_{2r}(p)),\mathsf{d}^{\delta})$, for $\iota$ as in Section 2.3 where $\mathsf{d}^{\delta}$ is the distance on $(N,g^{\delta})$. Since $|\nabla h(y)|_{g^{\delta}}\leq 2|\nabla f(y)|_{g}$ for any $y\in B_{2r}(p)$, we find
@@
where in the third inequality we applied a Poincaré inequality as in (2.4), recalling that on smooth manifolds it is possible to take the integral on the right hand side on the ball of the same radius that appears on the left hand side [SaloffCosteBook, Corollary 5.3.5]. ∎
What observed so far implies that $C^{0}$-Riemannian manifolds locally asymptotic at infinity to space forms are PI spaces, see the forthcoming Section 2.3. We recall their definition, leaving the interested reader to the seminal [HajlaszKoskela, Cheeger00] and to the survey [KleinerMackay].
[…]
@@
Let $R>0$ and fix $o\in M$. By assumptions and recalling Section 2.1, we can fix $\rho>R$ such that for any $x\in M\setminus B_{\rho}(o)$ there exists a diffeomorphism $F_{x}:(B_{3R}(x),g)\to(\mathbb{H}^{n}_{K},g_{K})$ that is $2$-biLipschitz with its image, where $g_{K}$ is the metric on $\mathbb{H}^{n}_{K}$, and whose image contains $B_{R}(F_{x}(x))$. Since $\mathbb{H}^{n}_{K}$ is PI, arguing as in the proof of Section 2.3 it follows that there exist $\overline{C}>0$ such that
@@
Next apply Section 2.3 with $p=o$ and $R=2\rho$. Hence for any $x\in M$ we have that either $x\in B_{2\rho}(o)$, or $x\in M\setminus B_{\rho}(o)$. Hence the fact that $(M,\mathsf{d},\mathcal{H}^{n})$ is PI follows putting together the previous inequalities with those given by Section 2.3. ∎
The next corollary states the local Hölder continuity of the isoperimetric profile of $C^{0}$-Riemannian manifolds that are $C^{0}$-locally asymptotic to a model of constant curvature. The proof essentially follows a classical path, see e.g. [ANP, Lemma 2.23]. However in this context we do not have an explicit asymptotic rate for the perimeter of balls of infinitesimal radii, which are often used to perturb competitors. In place of ball, we shall employ images of Euclidean balls through biLipschitz maps into the manifold, so to get a one-parameter increasing family of sets whose perimeter has an explicit rate as the parameter goes to zero.
[…]
@@
Fix $o\in M$. Since $M$ is $C^{0}_{\rm loc}$-asymptotic to $\mathbb{H}^{n}_{K}$, there exists $r>0$ such that for any $p\in M\setminus B_{r}(o)$ there exists a $2$-biLipschitz map $F_{p}$ from a Euclidean ball of sufficiently small radius to $M$ with $F_{p}(0)=p$. Combining with Section 2.3, we conclude that there exists $\overline{r}\in(0,1)$ such that for any $p\in M$ there exists a $2$-biLipschitz map $F_{p}:(B_{\overline{r}}^{\rm eu}(0),\mathsf{d}_{\rm eu})\to F_{p}(B_{\overline{r}}^{\rm eu}(0))\subset M$ such that $F_{p}(0)=p$. Hence we define the one-parameter family of sets $E_{t}(p)\vcentcolon=F_{p}(B_{t}^{\rm eu}(0))$, for any $p\in M$ and $t\in(0,\overline{r}]$. In particular, recalling also the representation of the perimeter Section 2.3 and the fact that the essential boundary is biLipschitz invariant,

### v3

@@ 1.4. The question of rigidity
In the smooth global PMT [67], one gets that the mass is zero if and only if the metric is flat. Restricting the attention to the smooth case, an easy combination of the local IMCF arguments leading to Theorem˜1.5 with the rigidity triggered by the constancy of the Hawking mass along the weak IMCF [38] allows us to prove the following quasi-local counterpart.
@@
for all the domains $\partial E_{t}\Subset\Omega$ evolving through the local IMCF, where $\mathfrak{m}_{H}(\partial E_{t})$ denotes the Hawking mass (3.25) of their boundary. We can then easily conclude invoking a rigidity argument by Huisken–Ilmanen [38]. We observe that a counterpart of Theorem˜1.7, with (1.5) replaced by
@@
has been already obtained by Mondino–Templeton-Browne by different means [55]. In fact, we stress that having proved the existence of a weak local IMCF allows us to derive a different proof of the result in [55]. Indeed, (1.7) would directly imply (1.6), thus allowing us to conclude using the rigidity argument of Huisken–Ilmanen as mentioned above, see also the proof of (1.5) below. Notice that the condition (1.5) makes sense also for $C^{0}$-metrics.
@@
Acknowledgments. A part of this research was carried out while G.A., M. F., and M.P. were visiting S.N. at UFABC. The joyful working conditions and the hospitality are gratefully acknowledged. Another part of the work was carried out while G. A. and M. F. were participating in the workshop “Recent Advances in Comparison Geometry” in Hangzhou. They thank the Banff International Research Station and the Institute of Advanced Studies in Mathematics for this enriching opportunity. G.A. is grateful to Paula Burkhardt-Guim for useful discussions around the topic of the paper. M. F. is grateful to Andrea Mondino for his interest and for having brought to his attention the paper [55]. M.P. is grateful to Roberto Ladu for clarifying discussions and stimulating advice. The authors are grateful to Luca Benatti, Luciano Mari, and Kai Xu for useful discussions around the topic of this paper. They are also grateful to Virginia Agostiniani, Alessandro Carlotto, Jeff Jauregui, Dan Lee, Lorenzo Mazzieri, and Christina Sormani for their interest in this work.
@@ 2.1. Basic definitions and properties of $C^{0}$-metrics
We briefly recall some metric properties of $C^{0}$-Riemannian manifolds $(M,g)$, and we refer the reader to [21, 62] for a more detailed discussion. As in the smooth case, the continuous Riemannian metric $g$ defines a length structure on absolutely continuous curves on $M$, which in turn gives raise to a distance $\mathsf{d}$ that induces the manifold topology. We always assume that $(M,\mathsf{d})$ is a complete metric space, so that by Hopf–Rinow Theorem the distance $\mathsf{d}$ is geodesic, and every closed ball is compact.
@@
[…]
@@
Remark 2.4.
In the notation of Section˜2.1, arguing as in the proof of Section˜2.1, it follows that for any $R,\varepsilon>0$ the map $F_{i}:(B_{R}(p),\mathsf{d})\to(F_{i}(B_{R}(p)),\mathsf{d}_{i})$ is $(1+\varepsilon)$-biLipschitz with its image for any $i$ large enough.
Lemma 2.5.
@@
[…]
@@ 2.2. BV functions and sets of finite perimeter
Let $(M,g)$ be a complete $C^{0}$-Riemannian manifold of dimension $n$. Then we can consider the complete and separable metric measure space $(M,\mathsf{d},\mathcal{H}^{n})$. Following [54, 2], we define the total variation $|Df|(B)\in[0,+\infty]$ of a function $f\in L^{1}_{\mathrm{loc}}(M)$ in a Borel set $B\subseteq M$ as
@@
then plugging ${\rm lip\,}f_{n}$ in place of $|\nabla f_{n}|$ in (2.1) yields the same quantity. Indeed, if $f\in\operatorname{Lip}_{\rm loc}(M)$, then ${\rm lip\,}f=|\nabla f|$ almost everywhere, see Appendix˜A.
@@
Given any $f\in\mathrm{Lip}_{\mathrm{loc}}(M)$, it holds that $|Df|$ is a locally finite measure and $|Df|={\rm lip\,}f\mathcal{H}^{n}=|\nabla f|\mathcal{H}^{n}$, see Appendix˜A. Moreover, we recall the coarea formula in our setting,
Theorem 2.7 (Coarea formula [54, Proposition 4.2]).
@@ 2.3. Technical Lemmas on $C^{0}$-Riemannian manifolds
In this section we prove several technical lemmas about $C^{0}$-Riemannian manifolds we are going to use throughout the paper. Notably, we prove: a precompactness theorem for $BV$ functions on converging sequences of continuous Riemannian manifolds (Section˜2.3); the fact that every continuous Riemannian manifold is PI on every ball (Section˜2.3); continuity of the isoperimetric profile of continuous Riemannian manifolds with $C^{0}$-controlled geometry at infinity (Section˜2.3); and a generalized existence theorem for the isoperimetric problem on continuous Riemannian manifolds with $C^{0}$-controlled geometry at infinity (Theorem˜2.16).
Lemma 2.9.
@@
For any $\varepsilon\in(0,1)$ let $g^{\varepsilon}$ be a smooth Riemannian metric on $\overline{B}_{20r}^{g}(p)$ such that $|g(v,v)-g^{\varepsilon}(v,v)|\leq\varepsilon g(v,v)$ for any $v\in T_{x}M$ and $x\in B_{20r}(p)$. Gluing the boundary of a smooth connected open domain $D$ such that $\overline{B}_{10r}^{g}(p)\subset D\subset\subset B_{20r}^{g}(p)$ with a half-cylinder $0,+\infty)\times\partial D$ and suitably extending the metric $g^{\varepsilon}$, we obtain a smooth complete Riemannian manifold $(N,g^{\varepsilon})$. Denote by $\mathsf{d}^{\varepsilon}$ the Riemannian distance on $(N,g^{\varepsilon})$ and by $\tilde{\mathsf{d}}$ the distance defined in Section˜2.1 with $F=\iota:B_{10r}^{g}(p)\to B_{10r}^{g}(p)\subset(N,g^{\varepsilon})$. Arguing as in the second part of the proof of Section˜2.1, it follows that $\tilde{\mathsf{d}}=\mathsf{d}^{\varepsilon}$ on $B_{r}^{g}(p)$. Indeed, as in the proof of Section˜2.1 we find that for any $x,y\in B^{g}_{r}(p)$ there exists a constant speed curve $\gamma:[0,1]\to B^{g}_{10r}(p)$ such that $\tilde{\mathsf{d}}(x,y)=\int_{0}^{1}|\gamma^{\prime}|_{g^{\varepsilon}}$, hence $\gamma$ is a geodesic for the metric $g^{\varepsilon}$. For every $x\in B^{g}_{r}(p)$, we have that for almost every $y\in B^{g}_{r}(p)$ there exists a unique, hence minimizing, geodesic in $(N,g^{\varepsilon})$ from $x$ to $y$. Then, for every $x\in B^{g}_{r}(p)$, we have that for almost every $y\in B^{g}_{r}(p)$ the curve $\gamma$ just obtained must be the minimizing geodesic from $x$ to $y$ in $(N,g^{\varepsilon})$. Hence, given $x\in B^{g}_{r}(p)$, there holds $\tilde{\mathsf{d}}(x,y)=\mathsf{d}^{\varepsilon}(x,y)$ for almost every $y\in B^{g}_{r}(p)$. Thus by continuity $\tilde{\mathsf{d}}=\mathsf{d}^{\varepsilon}$ on $B^{g}_{r}(p)$. Therefore choosing $\varepsilon$ small enough and eventually renaming $g^{\varepsilon}$ into $g^{\delta}$, the conclusion follows from Section˜2.1. ∎
Remark 2.10 (Representation of the perimeter).
@@
Indeed, fix on a compact ball $B\subset M$ a sequence $g_{i}$ of smooth metrics such that $g_{i}\to g$ uniformly on $B$ (see, e.g., Section˜2.3). Then, up to subsequences, $\partial^{e}E\cap B$ does not depend on the metrics $g_{i},g$ chosen in the definition (2.3) (compare also with item (1) in Section˜2.3). Moreover, by the very definition of Hausdorff measure, $(1-o_{i}(1))\mathcal{H}^{n-1}_{g_{i}}\llcorner B\leq\mathcal{H}^{n-1}_{g}\llcorner B\leq(1+o_{i}(1))\mathcal{H}^{n-1}_{g_{i}}\llcorner B$. From the classical De Giorgi-Federer’s theorem, $|D\chi_{E}|_{i}(\Omega)=\mathcal{H}^{n-1}_{g_{i}}(\partial^{e}E\cap\Omega)$ for every $\Omega\subset\subset B$. Thus, taking into account Appendix˜A one has that $|D\chi_{E}|_{i}(\Omega)\to|D\chi_{E}|(\Omega)$ for every Borel $\Omega\subset\subset B$, and thus we conclude $|D\chi_{E}|\llcorner B=\mathcal{H}^{n-1}\llcorner(\partial^{e}E\cap B)$. Since $B$ was arbitrary, we get the sought claim. In the following, if $E$ is a set of locally finite perimeter, integrals over its essential boundary are tacitly understood to be taken with respect to the perimeter measure.
Lemma 2.11.
Let $(M_{i},g_{i},p_{i})$ be a sequence of pointed $C^{0}$-Riemannian manifolds of dimension $n$ converging in $C^{0}$-sense to a pointed $C^{0}$-manifold $(M,g,p)$. Denote by $\mathsf{d},\mathsf{d}_{i}$ the Riemannian distances on $(M,g),(M_{i},g_{i})$ respectively. Let $f_{i}\in BV(M_{i})$ be such that $\sup_{i}\|f_{i}\|_{L^{1}(M_{i})}+|Df_{i}|(M_{i})<+\infty$. Then, up to subsequence, there exist $f\in BV(M)$, $R_{i}\nearrow+\infty$ and $(1+1/i)$-biLipschitz embeddings $F_{i}:(B_{R_{i}}(p),\mathsf{d})\to(M_{i},\mathsf{d}_{i})$ with $F_{i}(p)=p_{i}$ such that the functions $f_{i}\circ F_{i}$ converge to $f$ in $L^{1}_{\rm loc}(M)$. Moreover there holds
@@
By a diagonal argument, up to passing to a subsequence, by Section˜2.1 there exist $R_{i}\nearrow+\infty$ and $(1+1/i)$-biLipschitz embeddings $F_{i}:B_{R_{i}}(p)\to M_{i}$ with $F_{i}(p)=p_{i}$. Denote $h_{i}\vcentcolon=f_{i}\circ F_{i}$. For any $r>0$, for $i$ large we have that $h_{i}\in BV(B_{r}(p))$ with $\sup_{i}\|h_{i}\|_{L^{1}(B_{r}(p))}+|Dh_{i}|(B_{r}(p))<+\infty$. If we show that $h_{i}$ admits a subsequence converging in $L^{1}(B_{r/2}(p))$, the first part of the statement follows. Indeed, by Section˜2.3 we can find a smooth Riemannian manifold $(N,g^{\delta})$ such that it is well defined the inclusion $\iota:(B_{r}(p),\mathsf{d})\to(N,\mathsf{d}^{\delta})$ and $\iota$ is $2$-biLipschitz with its image. Hence $h_{i}$ can be seen as an equibounded sequence in $BV(B_{r}^{g}(p),g^{\delta})$. By classical precompactness we can extract a subsequence converging in $L^{1}(B_{r/2}^{g}(p),\mathcal{H}^{n}_{g^{\delta}})$. Since $\mathcal{H}^{n}_{g^{\delta}}$ and $\mathcal{H}^{n}_{g}$ are equivalent, the subsequence converges in $L^{1}(B_{r/2}(p))$ as well.
@@
The claims are well-known to hold true on manifolds with smooth Riemannian metrics as a consequence of the fact that one can always find a lower bound for the Ricci curvature on compact sets. In fact, on a smooth manifold, (1) is a consequence of the Bishop–Gromov comparison theorem [61, Lemma 7.1.4], (2) follows from [22, 63], and (3) is a consequence of (1) and (2) by [34, Theorem 9.7] and [3, Theorem 4.3, Remark 4.4]. Therefore claims (1) and (3) in the setting of the statement readily follow from Section˜2.3 applied with $r$ sufficiently large.
It remains to prove (2). Let $(N,g^{\delta})$ be given by Section˜2.3 with $\delta=1/4$ and radius equal to $3R$. Letting now $f\in\operatorname{Lip}_{\rm loc}(M)$, for $x\in B_{R}(p)$ and $r\leq R$, we have $h\vcentcolon=f\circ\iota^{-1}\in\operatorname{Lip}_{\rm loc}(\iota(B_{2r}(p)),\mathsf{d}^{\delta})$, for $\iota$ as in Section˜2.3 where $\mathsf{d}^{\delta}$ is the distance on $(N,g^{\delta})$. Since $|\nabla h(y)|_{g^{\delta}}\leq 2|\nabla f(y)|_{g}$ for any $y\in B_{2r}(p)$, we find
@@
where in the third inequality we applied a Poincaré inequality as in (2.4), recalling that on smooth manifolds it is possible to take the integral on the right hand side on the ball of the same radius that appears on the left hand side [66, Corollary 5.3.5]. ∎
What observed so far implies that $C^{0}$-Riemannian manifolds locally asymptotic at infinity to space forms are PI spaces, see the forthcoming Section˜2.3. We recall their definition, leaving the interested reader to the seminal [34, 24] and to the survey [45].
Definition 2.13 (PI space).
@@
Let $R>0$ and fix $o\in M$. By assumptions and recalling Section˜2.1, we can fix $\rho>R$ such that for any $x\in M\setminus B_{\rho}(o)$ there exists a diffeomorphism $F_{x}:(B_{3R}(x),g)\to(\mathbb{H}^{n}_{K},g_{K})$ that is $2$-biLipschitz with its image, where $g_{K}$ is the metric on $\mathbb{H}^{n}_{K}$, and whose image contains $B_{R}(F_{x}(x))$. Since $\mathbb{H}^{n}_{K}$ is PI, arguing as in the proof of Section˜2.3 it follows that there exist $\overline{C}>0$ such that
@@
Next apply Section˜2.3 with $p=o$ and $R=2\rho$. Hence for any $x\in M$ we have that either $x\in B_{2\rho}(o)$, or $x\in M\setminus B_{\rho}(o)$. Hence the fact that $(M,\mathsf{d},\mathcal{H}^{n})$ is PI follows putting together the previous inequalities with those given by Section˜2.3. ∎
The next corollary states the local Hölder continuity of the isoperimetric profile of $C^{0}$-Riemannian manifolds that are $C^{0}$-locally asymptotic to a model of constant curvature. The proof essentially follows a classical path, see e.g. [5, Lemma 2.23]. However in this context we do not have an explicit asymptotic rate for the perimeter of balls of infinitesimal radii, which are often used to perturb competitors. In place of ball, we shall employ images of Euclidean balls through biLipschitz maps into the manifold, so to get a one-parameter increasing family of sets whose perimeter has an explicit rate as the parameter goes to zero.
Corollary 2.15.
@@
Fix $o\in M$. Since $M$ is $C^{0}_{\rm loc}$-asymptotic to $\mathbb{H}^{n}_{K}$, there exists $r>0$ such that for any $p\in M\setminus B_{r}(o)$ there exists a $2$-biLipschitz map $F_{p}$ from a Euclidean ball of sufficiently small radius to $M$ with $F_{p}(0)=p$. Combining with Section˜2.3, we conclude that there exists $\overline{r}\in(0,1)$ such that for any $p\in M$ there exists a $2$-biLipschitz map $F_{p}:(B_{\overline{r}}^{\rm eu}(0),\mathsf{d}_{\rm eu})\to F_{p}(B_{\overline{r}}^{\rm eu}(0))\subset M$ such that $F_{p}(0)=p$. Hence we define the one-parameter family of sets $E_{t}(p)\vcentcolon=F_{p}(B_{t}^{\rm eu}(0))$, for any $p\in M$ and $t\in(0,\overline{r}]$. In particular, recalling also the representation of the perimeter Section˜2.3 and the fact that the essential boundary is biLipschitz invariant,